The partition coefficient of a base B (monoacidic) at a particular pH, in an octanol-water mixture
is 1.45. Consider an experiment where 3.5 g of B are shaken with 100 mL of octanol and 50 mL
of water.
Molar mass of B = 101.19 g mol-1
(i) Calculate the mass of B in the aqueous layer after shaking.
1
Expert's answer
2020-06-08T15:33:14-0400

By definition, 1.45 = [B]octanol / [B]water

number of moles B = 3.5/101.19 = 0.0346 moles

[B]octanol = (0.0346 - x)/0.100

[B]water = x/0.050

where x is number of moles B in water.

1.45 = (0.0346 - x)*0.050/0.100x = (0.0346 - x)/2x

x = 0.0089 moles B in water

mass = 0.0089*101.19 = 0.898 g B in water